# Pretrained Reversible Generation as Unsupervised Visual Representation Learning

### v3

@@ C.3 Details of Out-of-Distribution Experiments
Method Clean Average Corruption-$5$
@@
PGD [madry2017towards] $51.08$ $33.46$ $\downarrow 17.62$ $24.00$ $\downarrow 27.08$
PLAT [kireev2022effectiveness] $51.29$ $37.92$ $\downarrow 13.37$ $29.05$ $\downarrow 22.24$
@@
RSE [liu2018towards] $53.74$ $27.99$ $\downarrow 25.75$ $18.92$ $\downarrow 34.82$
ENResNet [wang2020enresnet] $49.26$ $25.83$ $\downarrow 23.43$ $19.01$ $\downarrow 30.25$
@@
AugMix [hendrycks2019augmix] $52.82$ $37.74$ $\downarrow 15.08$ $28.66$ $\downarrow 24.16$
AutoAug [cubuk2019autoaugment] $52.63$ $35.14$ $\downarrow 17.49$ $25.36$ $\downarrow 27.27$
@@
PDE+ [yuan2024pde] $53.72$ $39.41$ $\downarrow 14.31$ $30.32$ $\downarrow 23.40$
[…]
@@ C.4 The Number of Timesteps
C.5 Transfer Experiment on Dual-Task Datasets
Following the approach described in LABEL:sec:trans, we transfer the powerful text-to-image model SiT-XL [ma2024sit] from the community. However, instead of directly applying it to downstream classification tasks, we leverage its diffusion trajectory for feature extraction across multiple tasks. Specifically, we set the diffusion trajectory to 30 steps, extracting features from the 10-th reverse diffusion step for CIFAR classification and the 20-th step for Tiny-ImageNet.
Our algorithm, PRG-SiT-XL2, achieves an accuracy of 96.9% on CIFAR and 70.2% on Tiny-ImageNet.

### v6

@@ C.3 Details of Out-of-Distribution Experiments
Method Clean Average Corruption-55
@@
PGD [44] $51.08$ $33.46$ $\downarrow 17.62$ $24.00$ $\downarrow 27.08$
PLAT [32] $51.29$ $37.92$ $\downarrow 13.37$ $29.05$ $\downarrow 22.24$
@@
RSE [38] $53.74$ $27.99$ $\downarrow 25.75$ $18.92$ $\downarrow 34.82$
ENResNet [66] $49.26$ $25.83$ $\downarrow 23.43$ $19.01$ $\downarrow 30.25$
@@
AugMix [25] $52.82$ $37.74$ $\downarrow 15.08$ $28.66$ $\downarrow 24.16$
AutoAug [15] $52.63$ $35.14$ $\downarrow 17.49$ $25.36$ $\downarrow 27.27$
@@
PDE+ [71] $53.72$ $39.41$ $\downarrow 14.31$ $30.32$ $\downarrow 23.40$
PRG-ICFM-S (ours) $56.85$ $46.93$ $\downarrow 9.92$ $33.32$ $\downarrow 23.53$